SLOW-ELECTRON VELOCITY-MAP IMAGING OF NEGATIVE IONS: APPLICATIONS TO SPEC-TROSCOPY AND DYNAMICS

\underline{DANIEL NEUMARK}, Department of Chemistry, University of California at Berkeley, B64 Hildebrand Hall, Berkeley, CA 94720.

Slow electron velocity map imaging (SEVI) of negative ions is a high resolution variant of negative ion photoelectron spectroscopy that can be used with great effect to investigate spectroscopy and dynamics in radicals, clusters, and transition states. It is particularly powerful when combined with ion trapping and cooling. Several examples will be presented.